OpenStudy (anonymous):

Bobby ran 3/10 of a mile in 1/1/2 mintues. How many minutes per mile is this?

OpenStudy (anonymous):

1/1/2=3/2 (3/2)/(3/10) (3/2) *(10/3) time flip ((3*10)/(2*3)= 30/6 = 5 minute one mile